# ReSem3D: Refinable 3D Spatial Constraints via Fine-Grained Semantic Grounding for Generalizable Robotic Manipulation

### v3

@@ Abstract
Semantics-driven 3D spatial constraints align high-level semantic representations with low-level action spaces, facilitating the unification of task understanding and execution in robotic manipulation. The synergistic reasoning of Multimodal Large Language Models (MLLMs) and Vision Foundation Models (VFMs) enables cross-modal 3D spatial constraint construction. Nevertheless, existing methods have three key limitations: (1) coarse semantic granularity in constraint modeling, (2) lack of real-time closed-loop planning, (3) compromised robustness in semantically diverse environments. To address these challenges, we propose ReSem3D, a unified manipulation framework for semantically diverse environments, leveraging the synergy between VFMs and MLLMs to achieve fine-grained visual grounding and dynamically constructs hierarchical 3D spatial constraints for real-time manipulation. Specifically, the framework is driven by hierarchical recursive reasoning in MLLMs, which interact with VFMs to automatically construct 3D spatial constraints from natural language instructions and RGB-D observations in two stages: part-level extraction and region-level refinement. Subsequently, these constraints are encoded as real-time optimization objectives in joint space, enabling reactive behavior to dynamic disturbances. Extensive simulation and real-world experiments are conducted in semantically rich household and sparse chemical lab environments. The results demonstrate that ReSem3D performs diverse manipulation tasks under zero-shot conditions, exhibiting strong adaptability and generalization.
@@ I Introduction
With the advancement of robotic autonomy in unstructured environments, the growing diversity of semantic descriptions and the complexity of visual modeling have become critical challenges limiting system generalization and robustness [ref1]. Spatial constraints [ref2, ref3] serve as the critical interface between high-level semantic representations and low-level action spaces, enabling the mapping of abstract semantics into executable actions. Recent advances in cross-modal reasoning powered by Multimodal Large Language Models (MLLMs) [ref4] and Vision Foundation Models (VFMs) [ref5, ref6] have significantly improved visual grounding capability, which is crucial for constructing accurate 3D spatial constraints. For instance, tweezers-based grasping entails localizing the coarse semantic grasp region and reasoning about the 3D geometric relationships between the tweezer tips and the grasp point. Based on these inferred relations, spatial constraint models can be constructed and parsed in real-time by the constraint optimizer, thereby enabling closed-loop control and fine-grained manipulation in dynamic environments.
However, current methods lack an optimizable intermediate representation that effectively links linguistic intent with the action space, thereby hindering consistent alignment between semantic goals and physical execution. As a core intermediate representation, 3D spatial constraints parameterize the mapping from semantic intent to executable actions through a set of task-grounded poses. Nevertheless, the geometric reasoning and spatial localization capabilities of MLLMs remain insufficient for directly deriving fine-grained constraints from language. Prior efforts have leveraged MLLMs and VFMs to enhance language–vision semantic alignment [ref7, ref8], facilitating the construction of 3D spatial constraints for motion planning [ref9, ref10]. These constraints often remain limited to coarse localization and lack semantic-guided refinement. Furthermore, downstream controllers typically rely on task-space planning, which is limited by the stability and efficiency of inverse kinematics solvers. These limitations raise two critical challenges: 1) how to design a flexible and optimizable 3D spatial constraint model that automatically maps semantic intent to robotic actions, and 2) how to ensure stable closed-loop and real-time execution within the low-level action space in semantically diverse environments.
To overcome the first challenge, we propose a two-stage hierarchical 3D spatial constraint modeling approach driven by the synergistic reasoning of VFMs and MLLMs: part-level constraint extraction and region-level constraint refinement. The part-level extraction leverages VFMs to identify semantically relevant part-level regions from the RGB observation, while generating explicit digital markers as visual prompts to facilitate MLLMs in coarse-grained constraint extraction. The region-level refinement builds dense semantic grids within masked regions, enabling fine-grained semantic grounding of geometric features by MLLMs. Subsequently, 3D spatial constraints are encoded as cost functions parsed in real-time, and supplied to downstream optimizers. To tackle the second challenge, inspired by a finite-horizon optimal control approach grounded in a general-purpose physics simulator [ref11, ref12], we implement a real-time Model Predictive Path Integral (MPPI) control algorithm in the Isaac Gym platform [ref13], which delivers 15 Hz joint-space velocity commands instead of task-space commands for stable closed-loop execution.
@@
In addition, we develop a Task and Motion Planning (TAMP) framework that extends the hierarchical recursive prompting structure of Liang et al. [ref14], driven by MLLM-based reasoning, enabling autonomous and closed-loop manipulation from perception to action. The framework demonstrates semantics-driven adaptability by autonomously determining the required granularity of 3D spatial constraints from natural language instructions. It decomposes multi-stage tasks and generates precondition checks, postcondition evaluations, and optimizable cost functions necessary for the operation of each subtask. Building on this, a subtask-level backtracking strategy is formulated, leveraging precondition feasibility to initiate cross-stage replan, enabling behavior recovery and reactive control under dynamic perturbations.
@@
    We propose a two-stage hierarchical constraint modeling framework that integrates VFMs and MLLMs to build semantically adaptive, multi-granularity 3D spatial constraints, effectively bridging the gap between high-level semantics and low-level actions.
@@
    Closed-loop control for complex long-horizon tasks with fine-grained semantic grounding is achieved for the first time with the newly developed MLLM-driven TAMP framework, which performs autonomous multi-stage task decomposition with condition reasoning and cost optimization.
@@
    Through extensive simulations and real-world experiments in semantically rich household and sparse chemical lab environments, we successfully validate ReSem3D’s robust zero-shot capabilities, demonstrating strong generalization and reactive property.
@@ II Related Works
II-A Visual Grounding with Foundation Models
@@
Visual grounding refers to mapping semantic features from natural language to image regions and spatial entities, evolving from low-level perception to semantic spatial reasoning with the advancement of foundation models. Prior VFMs focused on robust image perception, leveraging self-supervised learning [ref15, ref16] and architectures including CNNs [ref17, ref18] and Vision Transformers [ref19], which produce visual encodings for downstream tasks like object detection [ref20, ref21], segmentation [ref22, ref23], and feature extraction [ref24]. Furthermore, pixel-level parsing was enhanced through latent feature encoding and embedded visual prompts [ref25, ref26]. Subsequently, Vision-Language Models (VLMs) perform multimodal semantic grounding through visual-linguistic representations, achieving notable progress in image-text retrieval [ref27] and open-vocabulary object detection [ref28], but still struggle with fine-grained visual grounding. Building on this, researchers introduced Large Language Models (LLMs) [ref29] into vision-language systems, leading to MLLMs [ref30] that incorporate visual features as context for multimodal reasoning. Despite this, MLLMs remain dependent on front-end visual encoders, limiting spatial precision and semantic understanding. Recent studies have explored the synergy between VFMs and MLLMs by embedding visual prompts for semantic alignment [ref7], but fine-grained visual grounding remains a critical limitation. To address this challenge, we propose a multi-granularity semantic understanding framework with hierarchical multimodal reasoning, featuring part-level extraction and region-level refinement to enable fine-grained visual grounding.
II-B Spatial Constraints in Robotic Manipulation
@@
Spatial constraints are essential for robotic manipulation, bridging high-level semantics and low-level actions to ensure task feasibility. Conventional methods construct constraints through explicit geometric modeling of object shapes, contact dynamics, and environmental structure [ref31, ref32] to enable motion planning. While physically interpretable and effective in structured scenarios, their dependence on accurate modeling constrains their flexibility in complex settings. Recently, perception-driven learning has been used to infer potential manipulation regions from visual observations, including affordance recognition [ref33], grasp pose detection [ref34, ref35], target pose estimation [ref36], and semantic keypoint generation [ref37]. These approaches have further extended to multimodal manipulation frameworks that incorporate foundation model reasoning [ref38, ref39]. While capable of adapting to object deformations and semantic changes, they still require manual data collection and struggle with fine-grained semantic parsing and precise modeling of complex geometric relationships. Current approaches integrate language and vision models to enable zero-shot semantic region perception [ref2, ref9]. Nevertheless, aforementioned methods typically lack fine-grained semantic localization and executable geometric specifications. To overcome these limitations, we propose a two-stage 3D spatial constraint modeling framework that hierarchically maps natural language instructions to semantic objectives, enabling fine-grained and executable spatial constraints.
@@ II-C Task and Motion Planning
Task and Motion Planning (TAMP) provides a pivotal framework connecting high-level task reasoning with low-level motion execution. Conventional approaches often leverage formal languages such as PDDL and HTN to model symbolic tasks, solving task sequences and motion trajectories through logic-geometric and mixed-integer programming [ref40, ref41]. Although offering strong interpretability, these methods rely on manually designed task models and action primitives, limiting their adaptability in open environments. With the advancement of LLMs, recent works have explored zero-shot task planning based on predefined motion primitives [ref42]. However, these approaches lack effective modeling of geometric constraints and environmental dynamics. Building on these developments, recent research has explored integrating LLMs with motion planning to enable high-level task decomposition and continuous action generation without predefined primitives [ref10]. Although such methods enhance the flexibility of TAMP, they often depend on online optimization in task space, posing challenges for real-time execution. In order to address these challenges, we employ a real-time MPPI optimizer built on Isaac Gym, leveraging joint velocity control to enhance reactive behavior during execution. Concurrently, subtask-level constraints and optimization objectives are automatically generated by the MLLM, enabling effective closed-loop control and zero-shot adaptation in dynamic environments.
Figure 1: Overall Framework. Given natural language instruction and RGB-D observations, VFM segments semantically relevant part-level regions and overlays visual prompts to facilitate initial constraint generation. Under the MLLM-driven TAMP framework, constraint modeling is conducted hierarchically in two stages: part-level extraction and region-level refinement. The resulting 3D spatial constraints are encoded as cost function for real-time parsing and solved in closed-loop by MPPI-based optimizer in Isaac Gym, enabling joint-space velocity control with point tracking.
@@ III Method
In this section, we first formulate ReSem3D as a real-time constrainted optimization problem to standardize task objectives and constraint. Subsequently, we propose a two-stage hierarchical spatial constraint model with semantic adaptivity for multi-granularity 3D reasoning. Thereafter, we derive a real-time closed-loop control strategy for solving the optimization problem. Finally, we integrate MLLM-driven reasoning to automate TAMP, enabling end-to-end closed-loop execution.
III-A ReSem3D Problem Formulation
As shown in Fig. 1, given a free-form natural language instruction $\mathcal{L}=\{L_{1},L_{2},\ldots,L_{n}\}$ from the user, $\mathcal{L}$ typically exhibits significant long-term dependencies and complex semantic reasoning. Directly generating complete 3D spatial constraints and optimizing them both compromises reasoning accuracy and complicates constraint modeling. Therefore, ReSem3D employs a hierarchical modeling strategy by decomposing the task into sequential subtasks and performing constraint reasoning and optimization progressively for each $L_{i}$, $i\in\{1,2,\ldots,n\}$.
Specifically, for the $i$-th subtask instruction $L_{i}$, the instantiation of the 3D spatial constraint model is defined as
@@ III-A ReSem3D Problem Formulation
where $O_{i}$ denotes the RGB-D sensor observations, $\mathcal{C}_{i}^{\text{init}}\in SE(3)$ denotes the initial set of spatial constraints mapped by $f$. In order to capture uncertainties in dynamic environments, we introduce external dynamic evolution and disturbances $\mathcal{E}_{i}$. Additionally, we define $\bm{T}_{e}(t)\in SE(3)$ as the end-effector pose at time $t$, and $S$ as the known kinematic and physical model. At each stage $i$, ReSem3D addresses task execution by resolving the constrained optimization problem in real-time:
@@
$\bm{v}(t),\ \bm{T}_{e}(t)\in\text{Feasible}\left(S\right)$
where $\mathcal{J}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ is the real-time optimized cost function, $g\left(\cdot\right)$ denotes the instantiation mapping that updates spatial constraints based on external disturbances, $\epsilon_{\text{pre}}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ and $\epsilon_{\text{post}}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ represent the precondition and postcondition constraints for subtask execution, with $\varepsilon_{\text{pre}}$ and $\varepsilon_{\text{post}}$ specifying their tolerance thresholds. The set $\text{Feasible}\left(S\right)$ enforces consistency with the robot kinematic and physical model, $\bm{v}(t)\in\mathbb{R}^{n}$ is the joint space velocity, serving as the decision variable, with $n$ denoting the degrees of freedom.
Given the previously described configurations, ReSem3D decomposes long-horizon manipulation tasks into multi-stage real-time constrained optimization problems, enabling closed-loop dynamic control for each sub-task. Subsequently, the processed decision variable $\bm{v}(t)$ directly serves as the joint velocity command, enabling the robot to perform tasks effectively in dynamic environments.
Figure 2: Part-Level Constraint Extraction: Original mask is generated via Fast Segment Anything (FastSAM), followed by mask filtering, part-level semantic consistency cluster, and centroid annotation to extract part-level spatial constraints within the MLLM-driven TAMP framework.
III-B Two-Stage Hierarchical 3D Spatial Constraint Model
To instantiate the mapping $f:(L_{i},O_{i})\rightarrow\mathcal{C}_{i}^{\text{init}}$, we leverage VFM to prompt the MLLM for fine-grained visual grounding, enabling the automated construction of a two-stage hierarchical 3D spatial constraint model comprising part-level constraint extraction and region-level constraint refinement.
III-B1 Part-Level Constraint Extraction
As illustrated in Fig. 2, given subtask instruction $L_{i}$ and RGB-D observation $O_{i}\in\mathbb{R}^{H\times W\times 4}$, we apply the pretrained image segmentation model [ref5] to produce the initial set of masks $M=\{m_{1},m_{2},\ldots,m_{g}\}$, where $m_{j}\in{\{0,\ 1\}}^{H\times W}$ and $\forall j\in\{1,\ldots,g\}$. Here, $H$ and $W$ denote the height and width of the image, respectively. To extract high-quality and discriminative part-level constraints, we design the following three-step post-processing strategy.

### v4

# ReSemAct: Advancing Fine-Grained Robotic Manipulation via Semantic Structuring and Affordance Refinement

@@ Abstract
Fine-grained robotic manipulation requires grounding natural language into appropriate affordance targets. However, most existing methods driven by foundation models often compress rich semantics into oversimplified affordances, preventing exploitation of implicit semantic information. To address these challenges, we present ReSemAct, a novel unified manipulation framework that introduces Semantic Structuring and Affordance Refinement (SSAR), powered by the automated synergistic reasoning between Multimodal Large Language Models (MLLMs) and Vision Foundation Models (VFMs). Specifically, the Semantic Structuring module derives a unified semantic affordance description from natural language and RGB observations, organizing affordance regions, implicit functional intent, and coarse affordance anchors into a structured representation for downstream refinement. Building upon this specification, the Affordance Refinement strategy instantiates two complementary flows that separately specialize geometry and position, yielding fine-grained affordance targets. These refined targets are then encoded as real-time joint-space optimization objectives, enabling reactive and robust manipulation in dynamic environments. Extensive simulation and real-world experiments are conducted in semantically rich household and sparse chemical lab environments. The results demonstrate that ReSemAct performs diverse tasks under zero-shot conditions, showcasing the robustness of SSAR with foundation models in fine-grained manipulation.
@@ I Introduction
With the advancement of robotic autonomy in unstructured environments, the growing diversity of semantic descriptions and the complexity of visual modeling have become critical challenges limiting system generalization and robustness [walke2023bridgedata]. Recent advances in cross-modal reasoning powered by Multimodal Large Language Models (MLLMs) [achiam2023gpt-4] and Vision Foundation Models (VFMs) [zhao2023fastsam] have significantly improved semantic grounding, which is crucial for constructing accurate affordance targets. For instance, tweezers-based grasping entails localizing the coarse semantic grasp region and reasoning about the geometric relationships between the tweezer tips and the grasp point. Based on these inferred relations, affordance targets can be constructed and parsed in real-time by the constraint optimizer, thereby enabling closed-loop control and fine-grained manipulation.
However, existing approaches typically map semantic information into coarse-grained affordance representations, lacking an optimizable intermediate structure that enables fine-grained semantic grounding [ju2025roboabc, li2025affgrasp]. Although Multimodal Large Language Models (MLLMs) have shown potential in cross-modal reasoning, their geometric inference and spatial localization capabilities remain insufficient for directly deriving fine-grained targets from natural language [li2024manipllm, liu2024moka]. Language-conditioned functional grasping [tang2025affordgrasp, tang2025foundationgrasp] provides partial associations between semantic intent and action execution, yet its expressiveness is confined to part-level granularity, limiting complex manipulation. Prior efforts have attempted to integrate MLLMs with VFMs to enhance language–vision semantic alignment [yang2023setofmark, yang2025magma] and support the automated construction of affordance targets, but these targets largely remain at coarse localization, lacking semantic-guided refinement [huang2023voxposer, huang2024rekep]. This highlights the urgent need for fine-grained semantic grounding to enable robotic manipulation. Inspired by these observations, we note that task-oriented grasping often implicitly encodes functional structures relevant to the scenario, while foundation models exhibit flexible semantic–visual conditioning. Based on this, we raise two key questions: 1) how to represent affordances with semantically flexible descriptions to support fine-grained grounding, and 2) how to develop a closed-loop framework that enables autonomous semantic grounding and execution in dynamic environments.
To address the first challenge, we propose Semantic Structuring and Affordance Refinement (SSAR). Specifically, the Semantic Structuring module derives a unified semantic affordance description by leveraging the synergistic reasoning of MLLMs and VFMs to extract affordance mask regions, implicit functional intent, and coarse affordance targets through multimodal reasoning, thereby forming structured semantic representation. Building upon this, the Affordance Refinement module instantiates geometric and positional flows to analyze the structured semantics affordance automatically, progressively refining the coarse affordance targets. Subsequently, the refined targets are encoded as optimization objectives within a Model Predictive Path Integral (MPPI) control [zhang2024mppi] running in real time on the Isaac Gym platform [makoviychuk2021isaacgym], with joint velocity optimization performed at 15 Hz.
@@
To address the second challenge, we develop a Task and Motion Planning (TAMP) framework. This framework extends the hierarchical recursive prompting structure of Liang et al. [liang2023cap] and is redeveloped under MLLM‑based reasoning, enabling the autonomous operation of SSAR and closed‑loop robotic execution. The framework automatically extracts and parses semantic information, supports multi‑stage task decomposition, and generates the necessary precondition checks, postcondition evaluations, and optimizable cost functions for each subtask. Furthermore, we propose a subtask‑level backtracking strategy that leverages precondition feasibility to trigger cross‑stage replanning, thereby achieving behavior recovery and reactive control under dynamic perturbations.
@@
    We propose Semantic Structuring and Affordance Refinement (SSAR), a new method that introduces the semantic affordance description through multimodal reasoning and the Affordance refinement flow to enable fine-grained semantic grounding.
@@
    Closed-loop control for diverse tasks with fine-grained semantic grounding is achieved with the newly developed MLLM-driven TAMP framework, which performs autonomous multi-stage task decomposition with condition reasoning and cost optimization.
@@
    Through extensive simulations and real-world experiments in semantically rich household and sparse chemical lab environments, we successfully validate ReSemAct’s semantic grounding capabilities, demonstrating strong generalization.
@@ II Related Works
II-A Semantic Grounding with Foundation Models
@@
Semantic grounding refers to mapping semantic features from natural language to image regions and spatial entities. Prior VFMs focused on image perception, leveraging self-supervised learning [chen2020cl, he2020momentumconstrast] and architectures including CNNs [lecun2002gradient, krizhevsky2017imagenet] and Vision Transformers [dosovitskiy2020image], which produce visual encodings for downstream tasks like object detection [girshick2014rich, redmon2016yolo], segmentation [long2015fullyconv, bad2017segnet], and feature extraction [touvron2021img]. Furthermore, pixel-level parsing was enhanced through latent feature encoding and embedded visual prompts [chen2022focalclick, zou2023samonce]. Subsequently, Vision-Language Models (VLMs) perform multimodal semantic grounding through visual-linguistic representations, achieving progress in image-text retrieval [radford2021learntrans] and open-vocabulary object detection [liu2024groundingdino], but still struggle with fine-grained semantic grounding. Building on this, researchers introduced Large Language Models (LLMs) [touvron2023llama] into vision-language systems, leading to MLLMs [liu2023visualins] that incorporate visual features as context for multimodal reasoning. Despite this, MLLMs remain dependent on front-end visual encoders, limiting spatial precision and semantic understanding. Recent studies have explored the synergy between VFMs and MLLMs by embedding visual prompts for semantic alignment [yang2023setofmark], but fine-grained semantic grounding remains a critical limitation. In contrast to the above methods, our work focuses on a structured semantic representation together with further refinement flows, framing fine-grained semantic grounding as an explicit and progressively optimized grounding strategy.
II-B Affordance-Based Robotic Manipulation
@@
Affordance are essential for robotic manipulation, bridging high-level semantics and low-level actions to ensure task feasibility. Conventional methods construct affordance through explicit geometric modeling of object shapes, contact dynamics, and environmental structure [karlbling2013tamp, kopicki2011manip] to enable motion planning. While physically interpretable and effective in structured scenarios, their dependence on accurate modeling constrains their flexibility in complex settings. Recently, perception-driven learning has been used to infer potential manipulation regions from visual observations, including affordance recognition [chu2019afford], grasp pose detection [fang2023anygrasp, pav2025grasp], target pose estimation [zheng2023hs-pose], and semantic keypoint generation [schmidt2016descriptor]. These approaches have further extended to multimodal manipulation frameworks that incorporate foundation model reasoning [tang2023graspgpt, li2025graspdex]. While capable of adapting to object deformations and semantic changes, they still require manual data collection and struggle with fine-grained semantic parsing and precise modeling of complex geometric relationships. Current approaches integrate language and vision models to enable zero-shot affordance region perception [huang2024copa, huang2024rekep]. Nevertheless, existing methods predominantly employ coarse affordance representations. By contrast, our work emphasizes semantic affordance descriptions that capture implicit intent, affordance regions, and affordance anchors, enabling the construction of semantically grounded affordance targets.
@@ II-C Task and Motion Planning
Task and Motion Planning (TAMP) provides a pivotal framework connecting high-level task reasoning with low-level motion execution. Conventional approaches often leverage formal languages such as PDDL and HTN to model symbolic tasks, solving task sequences and motion trajectories through logic-geometric and mixed-integer programming [dantam2016tamp, garrett2018tamp]. Although offering strong interpretability, these methods rely on manually designed task models and action primitives, limiting their adaptability in open environments. With the advancement of LLMs, recent works have explored zero-shot task planning based on predefined motion primitives [singh2023progprompt]. However, these approaches lack effective modeling of geometric constraints and environmental dynamics. Building on these developments, recent research has explored integrating LLMs with motion planning to enable high-level task decomposition and continuous action generation without predefined primitives [huang2023voxposer]. Although such methods enhance the flexibility of TAMP, they often depend on online optimization in task space, posing limitations for real-time execution. Our work implements an automated TAMP framework driven by MLLMs that couples high-level task reasoning with real-time joint-space constraint optimization, where subtask-level constraints and objectives are automatically generated to support closed-loop execution.
Figure 1: Overall Framework. Given natural language instructions and RGB-D observations, the synergistic reasoning between VFM and MLLM facilitates the semantic structuring representation. Within the MLLM-driven TAMP framework, the Semantic Structuring and Affordance Refinement Strategy performs geometric and positional flow optimization over the Semantic Affordance Description, thereby generating the refined affordance targets. These targets are then encoded as cost functions and solved in real-time by a closed-loop MPPI-based optimizer in Isaac Gym, enabling joint-space control with point tracking.
@@ III Method
In this section, we first formulate ReSemAct as a real-time constrained optimization problem with semantic grounding. Subsequently, we introduce Semantic Structuring and Affordance Refinement, which enables fine-grained semantic grounding through multimodal reasoning. Thereafter, we derive a real-time closed-loop control strategy for solving the optimization problem. Finally, we integrate MLLM-driven reasoning to automate TAMP, enabling closed-loop execution.
III-A ReSemAct Problem Formulation
As shown in Fig. 1, given a free-form natural language instruction $\mathcal{L}=\{L_{1},L_{2},\ldots,L_{n}\}$ from the user, $\mathcal{L}$ typically exhibits significant long-term dependencies and complex semantic reasoning. Simultaneous generation and optimization of affordance targets compromises reasoning accuracy and leads to overly coarse granularity. Therefore, ReSemAct employs a hierarchical modeling strategy by decomposing the task into sequential subtasks and progressively performs affordance reasoning and optimization for each $L_{i}$, $i\in\{1,2,\ldots,n\}$.
Specifically, for the $i$-th subtask instruction $L_{i}$, the instantiation of the semantic grounding is defined as
@@ III-A ReSemAct Problem Formulation
where $O_{i}$ denotes the RGB-D sensor observations, $\mathcal{C}_{i}^{\text{init}}\in SE(3)$ denotes the initial set of affordace targets mapped by $f$. In order to capture uncertainties in dynamic environments, we introduce external dynamic evolution and disturbances $\mathcal{E}_{i}$. Additionally, we define $\bm{T}_{e}(t)\in SE(3)$ as the end-effector pose at time $t$, and $S$ as the known kinematic and physical model. At each stage $i$, ReSemAct addresses task execution by resolving the constrained optimization problem in real-time:
@@
$\bm{v}(t),\ \bm{T}_{e}(t)\in\text{Feasible}(S)$
where $\mathcal{J}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ is the real-time optimized cost function, $g\left(\cdot\right)$ denotes the instantiation mapping that updates affordance targets based on external disturbances, $\epsilon_{\text{pre}}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ and $\epsilon_{\text{post}}\left(\bm{T}_{e}(t),\mathcal{C}_{i}\right)$ represent the precondition and postcondition constraints for subtask execution, with $\varepsilon_{\text{pre}}$ and $\varepsilon_{\text{post}}$ specifying their tolerance thresholds. The set $\text{Feasible}\left(S\right)$ enforces consistency with the robot kinematic and physical model, $\bm{v}(t)\in\mathbb{R}^{n}$ is the joint space velocity, serving as the decision variable, with $n$ denoting the degrees of freedom.
Building on the above configurations, ReSemAct decomposes long-horizon manipulation tasks into multi-stage real-time constrained optimization problems, achieving semantic grounding and closed-loop dynamic control for each sub-task. Subsequently, the processed decision variable $\bm{v}(t)$ directly serves as the joint velocity command, enabling the robot to perform tasks effectively in dynamic environments.
Figure 2: Semantic Structuring Representation. By extracting the implicit intent of instructions, the affordance regions from scene masks, and the coarse-grained affordance anchors of objects, the semantic affordance description is constructed within the MLLM-driven TAMP framework.
III-B Semantic Structuring and Affordance Refinement
To instantiate $f:(L_{i},O_{i})\rightarrow\mathcal{C}_{i}^{\text{init}}$, we leverages multimodal reasoning to produce the semantic affordance description, which is transformed via geometric and positional refinement flows to obtain fine-grained semantic grounding.
III-B1 Semantic Structuring Representation
As illustrated in Fig. 2, we define a semantic affordance description to construct an optimizable and interpretable semantic structure for downstream target refinement: the affordance region indicates the interactable space, the implicit functional intent specifies the fine-grained requirements, and the coarse affordance anchor provides the initial target for optimization.
Implicit Semantic Intent. To effectively leverage the implicit intent embedded in the given subtask instruction $L_{i}$, we extract its geometric and positional cues and reformulate them as an implicit functional intent $S_{\text{intent}}$ through recursive reasoning within the MLLM-driven TAMP framework.
Affordance Region. Given RGB-D observation $O_{i}\in\mathbb{R}^{H\times W\times 4}$, we apply the pretrained image segmentation model [zhao2023fastsam] to produce the initial set of masks $M=\{m_{1},m_{2},\ldots,m_{g}\}$, where $m_{j}\in{\{0,\ 1\}}^{H\times W}$ and $\forall j\in\{1,\ldots,g\}$. Here, $H$ and $W$ denote the height and width of the image, respectively. To extract the affordance regions $M_{\text{region}}$ of objects, the following two steps are designed: